OpenStudy (anonymous):

Jose is paying for a two dollar drink using bills in his wallet. He has four one dollar bills, two five dollar bills, and two ten dollar bills. If he selects two bills at random, one at a time from his wallet, without replacement, what is the probability that he will choose two one dollar bills to pay for the drink? Show your work.

OpenStudy (anonymous):

whoa thats alot of money in problem

OpenStudy (anonymous):

what if i dont wanna show my work :O

OpenStudy (anonymous):

go away then

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

NO you cant make me, im adorable remember? :D and i sent you a message, you never replied

OpenStudy (anonymous):

The fact that the outcome of choosing the first bill affects the outcome of choosing the second bill makes these events dependend. ( without replacement) E1 = Event that the first bill is a one dollar bill. E2 = Event that the second bill is a one dollar bill. P(E1) and E2) = P(E1) and P(E2|E1) P(E1) and P(E2|E1) = P(E1)*P(E2|E1) There are 4 + 2 + 2 = 8 bills to choose from. a) P(E1) = 4/8 prob. of choosing 1 of the 4 one dollar bills from 8 bills b) P(E2|E1) = 3/7 prob. of choosing 1 of the remaining 3 one dollar bills from remaining 7 bills c) P(E1 and E2) = P(E1)*P(E2|E1) =(4/8) * (3/7) = 3/14 Answer is 3/14 = 0.214286